Hello:
I really need your help. I have tried all possible ways to solve this equation but I am confused on how to start to solve it. Here is the question:
Find the point of intersection of the given straight lines.
x - 4y = -2
x + 2y = 4
Please help me with this problem. Thank you.

Found 2 solutions by rapaljer, sabanasir:
Answer by rapaljer(4671)

There are several ways to do this. The elimination method usually works well when the equations are both in standard form, as in this case.
x - 4y = -2
x + 2y = 4

In order to eliminate the x variable, multiply the first equation by -1, and rewrite the second equation as it is:
-x + 4y = 2
x + 2y = 4

Now add these two equations together giving you
6y = 6
y=1

Substitute the y = 1 back into either equation, say the first one:
x - 4y = -2
x - 4(1) = -2
x-4 = -2

Add 4 to each side:
x-4+4 = -2 + 4
x=2

Check in the second equation, by letting x= 2 and y = 1:
x + 2y = 4
2 + 2(1) = 4
It checks!

R^2 at SCC


Answer by sabanasir(22)

hmm it goes this way....
the idea first is to get a combined equation of the two either by adding them or subtracting them like...:
i can easily subtract the two equations:
(1) x-4y=-2
(2) x+2y=4
subtracting will get us free from one variable 'x' cancelled out.
+x-4y=-2
+x+2y=4
(-)
__________
-6y=-6
6y=6
y=1
okkkk now v get the value of y as 1.
we cud put y's value in either of the equation to get the value of 'x'
e.g i put it in (1)
x-4y=-2
put y=1
x-4(1)=-2
x-4=-2
x=-2+4
x=2
so the intersecting points are (x,y)=(2,1)